(a) The heat flux through the faces at the ends of bar is found to be proportional to un au/an at the ends. If the bar is perfectly insulated, also at the ends x 0 and x L are adiabatic conditions, Q1 ux(0, t) = 0 0 (2'7)*n prove that the solution of the heat transfer problem above (adiabatic conditions at both ends) gives as, 2 an: nnx u(x, t) Ao t An cos n-1 where Ao and An are an arbitrary constant The heat equation is: au at ax2 (17 marks) (b) If L and a 1 for the solution of heat transfer problem in Q1 (a), find the temperature in the bar with the initial temperature, f(x)= k = constant. (3 marks)
A half-range expansions given as the following function Q2 2k for 0
Answer #1

Heat of in ng By variable sepob (47) ucx t)=XTH (lat) Retg value in 0 XT X (Pos) t+xY 39 (xx -duit T= 3e ut Grorurxta hinan)given initial tepere tue ftw) = K(CONtont) fra Se en d L- TH)= C3ett K= C3 e 72 Constant -At This in khp. wf L-Tgiven initial tepere tue ftw) = K(CONtont) fra Se en d L- TH)= C3ett K= C3 e 72 Constant -At This in khp. wf L-T